OpenStudy (anonymous):

Find each percent of change and round to the nearest tenth. 1. 10 to 25 2. 70 to 82

OpenStudy (anonymous):

Percent change = new - old / old

OpenStudy (anonymous):

so for the first one you have 25 as your new, 10 as your old

OpenStudy (anonymous):

So would it be 25-10=15 divided by 10

OpenStudy (anonymous):

so it'd be \[\frac{25-10}{10}=\frac{15}{10}=1.5 x 100 = 150\]

OpenStudy (anonymous):

150% is your first

OpenStudy (anonymous):

the second should be easy since you did the first right

OpenStudy (anonymous):

Uhhh ok...

OpenStudy (anonymous):

answer should be 20%

OpenStudy (anonymous):

for the second one???